# Towards Large-Scale In-Context Reinforcement Learning by Meta-Training in Randomized Worlds

### v2

@@ C.1 Data Synthesis
Table 4: Summarizing the data synthesis strategies of different ICRL methods.
@@
OmniRL (Ours) $\mathfrak{o}$, $\mathfrak{q}$, $\mathfrak{m}$, $\mathfrak{r}$, $\mathfrak{o}^{\varepsilon}$ $\mathfrak{o}$
The data synthesis pipeline of OmniRL involves generating diverse trajectories $h$ using a variety of behavior policies and creating step-wise labels $l$ with an oracle policy $\pi^{\mathfrak{o}}$. This pipeline is detailed in Algorithm 2. We incorporate at least five distinct types of agents:
@@
With these notations, Table 4 can be used to represent not only the data synthesis pipeline of OmniRL but also the previous imitation meta-training-based ICRL methods, including AD, ADε, and DPT, as shown in Table 4. Notably, the synthesis pipeline of OmniRL is most similar to that of DPT. However, there are key differences: OmniRL employs a more diverse set of behavior policies and incorporates step-wise supervision (SS).
@@ 6 $\mathfrak{r}$ randomized policy (including perturbed action in $\mathfrak{o}^{\varepsilon}$)
For the prior information $g$, we assign eight different IDs to the actions with $g\in[0,7]$ which originated from $8$ types of different agents, as shown in Table 5. Specifically, we exclude the actions generated by the seven types of different agents and reserve $g=7$. This reserved ID is used to replace the action ID approximately $15\%$ of the time steps with the data synthesis pipeline of Algorithm 2.
@@
Addressing terminal states: The presence of terminal and truncation states necessitates special handling in reinforcement learning. In OmniRL, we avoid explicitly adding a terminal or truncation token to the sequence. Instead, we encode terminal and truncation states by introducing an additional action $a$, which is maintained as distinct from the standard action space $\mathcal{A}$. Additionally, we assign a reward of 0 and set the prior information $p=7$ for these special steps, as depicted in Figure 13.
@@ C.2 Meta-Training Details
Model structures: Before injection into causal models, the states ($s_{t}$) and actions ($a_{t}$) are encoded using embedding layers with a hidden size of $512$. The rewards ($r_{t}$) are treated as continuous features encoded by $1\times 512$ linear layer. The sequence model has a hidden size of $512$, inner hidden size of $1024$, hidden ratio of $2$, and block number of $18$ for RWKV-7. The model has approximately 43.6M total parameters (42.9M in RWKV-7 blocks), as shown in Table 6. We employ the open-source implementation of flash-linear-attention https://github.com/fla-org/flash-linear-attention.
@@
           forward: update $\phi_{k-1}\rightarrow\phi_{k}$ based on Equation 5, $\phi_{k-1}$, $h_{T_{k}:T_{k+1}}$ and $l_{T_{k}:T_{k+1}}$
@@
Meta-training: Algorithm 3 outlines the detailed process of the meta-training procedure. Notably, we perform the backward pass segment-wise and accumulate the gradients. The gradients are not applied until the end of a sequence. We utilize a constant segment length $T_{k+1}-T_{k}=2K$, which results in six backward passes for $T=12K$ before applying the gradient.
Table 7 provides an overview of the primary datasets used in this study. For the $\mathcal{D}_{\text{Large}}$ dataset, the state space size $n_{s}$ is uniformly sampled from the range $[16,128]$ to ensure robustness across varying state spaces. To evaluate the extrapolation capability of the model trained on $\mathcal{D}_{\text{Large}}$, we conducted a validation test with a context length of 1 million steps and observed that the loss began to gradually increase beyond $80K$ steps. Building upon this observation, we incorporated a post-training stage for long sequences with a context length of $512K$, the dataset is denoted as $\mathcal{D}_{\text{Long}}$.
@@
           sample: $a_{t}\sim p^{\theta}_{t}$ with Equation 3
@@
agent $\text{agent reward}=\begin{cases}1,&\text{if reach goal}\\ 0.08,&\text{if distance to goal decrease}\\ -0.12,&\text{if distance to goal increase }\\ -0.04,&\text{if still }\\ 0,&\text{if finish}\end{cases}$
shared $\text{shared reward}=\sum_{i=1}^{2}\text{agent reward}_{i}$
@@ C.3 Evaluation Details
As shown in Algorithm 4, since the episode length and baseline average episodic reward vary significantly across different tasks, we normalize the episodic reward using the oracle policy ($\mathfrak{o}$) and the uniform random policy ($\mathfrak{r}$). This normalization represents the percentage of oracle performance achieved. For AnyMDP, the evaluation averages the performances over $64$ variant unseen tasks. For Gymnasium tasks, the evaluation is conducted by averaging the results over $3$ runs on the same task.
@@
We apply some reward shaping to Gymnasium tasks as shown in Figure 14. OmniRL supports $n_{s}\leq 128$ and $n_{a}\leq 5$. For environments with $n_{a}<5$, we find directly setting $a=a\mod n_{a}$ is enough, which also demonstrates the generalizability of OmniRL to variant action spaces.
@@ Appendix D Additional Empirical Results
D.1 OmniRL validations
@@
OmniRL achieves automatic trade-off between exploration and exploitation. Previous studies have noted that in-context reinforcement learning (ICRL) can automatically balance exploration and exploitation. This phenomenon has been theoretically linked to posterior sampling. In Figure 15, we illustrate the entropy of the decision-making process as a function of steps within the context. When compared to Figure 3, we observe that the decrease in loss ($\mathcal{L}_{t}$) is primarily driven by the reduction in the entropy of the policy. Specifically, the agent initially assigns equal probabilities to all actions, reflecting an exploratory phase. As more contextual information accumulates, the agent gradually converges its choices, thereby transitioning towards exploitation. This empirical finding suggests that imitating an optimal policy (oracle) is sufficient to achieve an automatic balance between exploration and exploitation.
D.2 Additional Evaluation on Gymnasium
Figure 16 and Figure 17 demonstrate OmniRL’s online-RL, offline-RL, and imitation learning capabilities toward diverse unseen tasks.
@@ D.2 Additional Evaluation on Gymnasium
D.3 Can mainstream LLMs do ICRL?
{mdframed}
@@ D.3 Can mainstream LLMs do ICRL?
[roundcorner=10pt] You are playing the Frozen Lake game. The environment is a 4x4 grid where you need to maximize the success rate by reaching the goal (+1) without falling into holes (-1). You can move in four directions: left, down, right, and up (represented as 0, 1, 2, 3 respectively). You will receive the current state and need to provide the optimal action based on your learning. When asked for the optimal action, your response must be an integer ranging from 0 to 3, and no other context is permitted. There are two kind of request types:
1.integer: the integer is the current state, and you need to provide the optimal action.
2.list: The list contains one or more tuples, where each tuple contains the last state, action taken, reward received, and next state. To save time, you don’t need to respond when receiving a list.
You will play the game multiple times. A game ends when the reward is -1 or 1, try to get a higher success rate.
Note: I am asking you to play this game, not to find a coding solution or method.
You will be provided with a conversation history. The latest prompt is the current state, and others are the list of sequential environment feedback history in tuple type. Each tuple contains four values, the first one is state, the second one is action, the third one is reward and the fourth one is next state.
Your response must be an integer from 0 to 3 during the entire chat.
If you find the last state is equal to the next state, your policy in the last state can’t be this action.
If you find the reward in the tuple is -1, your policy in the last state can’t be this action.
You need to get to the goal as soon as possible.
Figure 18: Prompts for LLM to initialize the Lake4×$\times 4$ 4 (Slippery) task without a global map
{mdframed}
[roundcorner=10pt] There is a game with the following basic description and rules:
Frozen Lake involves crossing a frozen lake from the start to the goal without falling into any holes by walking over the frozen lake. The player may not always move in the intended direction due to the slippery nature of the frozen lake.
The game starts with the player at location [0,0] of the frozen lake grid world, with the goal located at the far extent of the world, for example, [3,3] for the 4x4 environment.
Holes in the ice are distributed in set locations when using a pre-determined map or in random locations when a random map is generated.
The player makes moves until they reach the goal or fall into a hole.
The lake is slippery, so the player may move perpendicular to the intended direction sometimes.
If the intended direction is to the left, the actual move may be to the left, up, or down, with the corresponding probability distribution: P(move left) = 1/3, P(move up) = 1/3, P(move down) = 1/3. If the intended direction is to the right, the actual move may be to the right, up, or down, with the corresponding probability distribution: P(move right) = 1/3, P(move up) = 1/3, P(move down) = 1/3. If the intended direction is up, the actual move may be up, left, or right, with the corresponding probability distribution: P(move up) = 1/3, P(move left) = 1/3, P(move right) = 1/3. If the intended direction is down, the actual move may be down, left, or right, with the corresponding probability distribution: P(move down) = 1/3, P(move left) = 1/3, P(move right) = 1/3. You are given a 4x4 map where:
S represents the start.
F represents the frozen surface that can be walked on.
H represents a hole; falling into it will return the player to the start.
G represents the goal.
The map is as follows:
The first row from left to right is "SFFF".
The second row from left to right is "FHFH".
The third row from left to right is "FFFH".
The fourth row from left to right is "HFFG".
Please determine the optimal policy that maximizes the success rate of safely reaching the goal from the start. The optimal policy is the intended direction at each map location, where actions 0, 1, 2, and 3 represent moving left, down, right, and up, respectively.
Note: You are not required to write code to solve this problem; instead, directly provide the optimal policy.
Figure 19: Prompts for LLM to solve Lake4×$\times 4$ 4 (Slippery) with global map
{mdframed}
@@
[roundcorner=10pt] In this particular environment, you should focus more on strategies to avoid encountering pitfalls rather than concentrating on reaching the goal. A reevaluation of the decision-making process in this context may yield more effective outcomes.
Figure 20: Additional guide (hint) for LLM to solve Lake4×$\times 4$ 4 (Slippery) with global map
{mdframed}
[roundcorner=10pt] The optimal policy for each state (row, column) is as follows, where each cell lists the intended direction (action) that maximizes the probability of reaching the goal:
- **(0,0)**: Down (1) - **(0,1)**: Down (1) - **(0,2)**: Down (1) - **(0,3)**: Up (3) - **(1,0)**: Down (1) - **(1,2)**: Down (1) - **(2,0)**: Right (2) - **(2,1)**: Down (1) - **(2,2)**: Down (1) - **(3,1)**: Right (2) - **(3,2)**: Right (2)
@@
**Explanation**: - **Top row (0,0 to 0,3)**: Prioritize moving down to avoid the risky rightward path toward holes. - **Middle rows (1,0 and 1,2)**: Move down to progress toward safer lower rows. - **Lower rows (2,0 to 3,2)**: Navigate rightward toward the goal, avoiding holes. For example, (3,1) and (3,2) move right to reach the goal directly. - **(0,3)**: Move up (intended) to bias movement leftward (safer) and avoid falling into the hole at (1,3).
Figure 21: Responses from DeepSeek-R1 for Lake4×$\times 4$ 4 (Slippery) with global map
{mdframed}
[roundcorner=10pt] The optimal policy, revised to prioritize avoiding holes, is as follows:
- **(0,0)**: Down (1) *Avoids (0,1)’s risks by favoring transitions to (1,0) or (0,1) without direct hole exposure.*
- **(0,1)**: Right (2) *Moves toward (0,2) or (1,1) (hole). While risky, all actions here have 1/3 hole risk; right progresses toward safer (0,2).*
- **(0,2)**: Right (2) *Directs to (0,3) instead of down to (1,2), which is high-risk.*
- **(0,3)**: Up (3) *Biases movement left (to (0,2)) and avoids the hole at (1,3).*
- **(1,0)**: Left (0) *Avoids 1/3 chance of moving right into (1,1) (hole). Transitions to (0,0) or (2,0) safely.*
- **(1,2)**: Right (2) *Despite 1/3 hole risk, it allows progress to (2,2) or (0,2). Left (action 0) has equal risk but less progress.*
- **(2,0)**: Up (3) *Avoids 1/3 chance of moving down into (3,0) (hole). Routes to (1,0) or (2,1).*
- **(2,1)**: Down (1) *Safely moves to (3,1) (frozen), progressing toward the goal.*
- **(2,2)**: Down (1) *Reaches (3,2), a safe step toward the goal.*
- **(3,1)**: Right (2) *Directly progresses to (3,2), then (3,3) (goal).*
- **(3,2)**: Right (2) *Final step to the goal (3,3).*
### Key Adjustments from Original Policy: 1. **(1,0)**: Changed to **left** (0) to avoid 1/3 chance of falling into (1,1). 2. **(1,2)**: Changed to **right** (2) instead of down, reducing hole risk from 2/3 to 1/3. 3. **(2,0)**: Changed to **up** (3) to avoid 1/3 chance of entering (3,0). 4. **(0,1)**: **Right** (2) minimizes repeated exposure to (1,1) compared to down.
### Path Safety Focus: - Avoids actions with >1/3 hole probability (e.g., down from (1,2)). - Prioritizes paths with lower immediate risk, even if longer (e.g., routing through (2,0) → (2,1) → (3,1) → goal).
Figure 22: Responses from DeepSeek-R1 for Lake4×$\times 4$ 4 (Slippery) with global map and hint
@@
We also investigate whether a well-pretrained LLM can naturally solve the mentioned. To circumvent the lack of common sense in AnyMDP tasks, we primarily conducted tests in the FrozenLake task with DeepSeek-R1[76] in two modes:
@@
    Similar to the evaluation of standard ICRL, we do not provide the agent with the map. Instead, we report only the state ID and reward of the agent. The initial prompts used to initiate the evaluation are shown in Figure 18.
@@
    We initially provide the global map to the $DeepSeek-R1$ and then commence the interaction. In this mode, the LLM can leverage the global map to make decisions. The prompts are shown in Figure 19 and Figure 20
As shown in Table 8 and Figure 23 (results excecuted by following the responses of DeepSeek-R1 (version 2025/03) in Figure 21 and Figure 22), LLM agents are only able to solve the FrozenLake (non-slippery) environment when provided with a global map. Without access to a global map, we conducted extensive interactions between LLM agents and the environment, running up to $500$ episodes ($100,000$ steps). Despite these efforts, the agents failed to solve even the non-slippery variant of the task, achieving scores that were comparable to those of a random policy.

### v3

@@ C.1 Data Synthesis
Table 4: Summarizing the data synthesis strategies of different methods under the Decoupled Policy Distillation (DPD) framework.
@@
Decoupled Policy Distillation (Ours) $\mathfrak{o}$, $\mathfrak{q}$, $\mathfrak{m}$, $\mathfrak{r}$, $\mathfrak{o}^{\varepsilon}$ $\mathfrak{o}$
The data synthesis pipeline of OmniRL involves generating diverse trajectories $h$ using a variety of behavior policies and creating step-wise labels $l$ with an oracle policy $\pi^{\mathfrak{o}}$. This pipeline is detailed in Algorithm˜2. We incorporate at least five distinct types of agents:
@@
With these notations, Table˜4 can be used to represent not only the data synthesis pipeline of OmniRL but also the previous imitation meta-training-based ICRL methods, including AD, ADε, and DPT, as shown in Table˜4. Notably, the synthesis pipeline of OmniRL is most similar to that of DPT. However, there are key differences: OmniRL employs a more diverse set of behavior policies and incorporates step-wise supervision (SS).
@@ 6 $\mathfrak{r}$ randomized policy (including perturbed action in $\mathfrak{o}^{\varepsilon}$)
For the prior information $g$, we assign eight different IDs to the actions with $g\in[0,7]$ which originated from $8$ types of different agents, as shown in Table˜5. Specifically, we exclude the actions generated by the seven types of different agents and reserve $g=7$. This reserved ID is used to replace the action ID approximately $15\%$ of the time steps with the data synthesis pipeline of Algorithm˜2.
@@
Addressing terminal states: The presence of terminal and truncation states necessitates special handling in reinforcement learning. In OmniRL, we avoid explicitly adding a terminal or truncation token to the sequence. Instead, we encode terminal and truncation states by introducing an additional action $a$, which is maintained as distinct from the standard action space $\mathcal{A}$. Additionally, we assign a reward of 0 and set the prior information $p=7$ for these special steps, as depicted in Figure˜12.
@@ C.2 Meta-Training Details
Model structures: Before injection into causal models, the states ($s_{t}$) and actions ($a_{t}$) are encoded using embedding layers with a hidden size of $512$. The rewards ($r_{t}$) are treated as continuous features encoded by $1\times 512$ linear layer. The sequence model has a hidden size of $512$, inner hidden size of $1024$, hidden ratio of $2$, and block number of $18$ for RWKV-7. The model has approximately 43.6M total parameters (42.9M in RWKV-7 blocks), as shown in Table˜6. We employ the open-source implementation of flash-linear-attention $\spadesuit$$\spadesuit\spadesuithttps$://github.com/fla-org/flash-linear-attention.
@@
    forward: update $\phi_{k-1}\rightarrow\phi_{k}$ based on Equation˜4, $\phi_{k-1}$, $h_{T_{k}:T_{k+1}}$ and $l_{T_{k}:T_{k+1}}$
@@
Meta-training: Algorithm˜3 outlines the detailed process of the meta-training procedure. Notably, we perform the backward pass segment-wise and accumulate the gradients. The gradients are not applied until the end of a sequence. We utilize a constant segment length $T_{k+1}-T_{k}=2K$, which results in six backward passes for $T=12K$ before applying the gradient.
Table˜7 provides an overview of the primary datasets used in this study. For the $\mathcal{D}_{\text{Large}}$ dataset, the state space size $n_{s}$ is uniformly sampled from the range $[16,128]$ to ensure robustness across varying state spaces. To evaluate the extrapolation capability of the model trained on $\mathcal{D}_{\text{Large}}$, we conducted a validation test with a context length of 1 million steps and observed that the loss began to gradually increase beyond $80K$ steps. Building upon this observation, we incorporated a post-training stage for long sequences with a context length of $512K$, the dataset is denoted as $\mathcal{D}_{\text{Long}}$.
@@
    sample: $a_{t}\sim p^{\theta}_{t}$ with Equation˜4
@@
[…]
$\text{shared reward}=\sum_{i=1}^{2}\text{agent reward}_{i}$
@@ C.3 Evaluation Details
As shown in Algorithm˜4, since the episode length and baseline average episodic reward vary significantly across different tasks, we normalize the episodic reward using the oracle policy ($\mathfrak{o}$) and the uniform random policy ($\mathfrak{r}$). This normalization represents the percentage of oracle performance achieved. For AnyMDP, the evaluation averages the performances over $64$ variant unseen tasks. For Gymnasium tasks, the evaluation is conducted by averaging the results over $3$ runs on the same task.
@@
We apply some reward shaping to Gymnasium tasks as shown in Figure˜13. OmniRL supports $n_{s}\leq 128$ and $n_{a}\leq 5$. For environments with $n_{a}<5$, we find directly setting $a=a\mod n_{a}$ is enough, which also demonstrates the generalizability of OmniRL to variant action spaces.
@@ Appendix D Additional Empirical Results
D.1 AnyMDP as a long-context benchmark for procedural memory
Figure 14: Comparison of meta-training dynamics across AnyMDP dataset of $6B$ time steps and step-wise loss (lower is better) on the validation set $\mathcal{D}(\mathcal{T}_{tst}(n_{s}\in[16,128],n_{a}=5))$ for different linear-attention models.
Figure Figure˜14 reports the performance of different linear-attention models throughout AnyMDP training and on the held-out AnyMDP evaluation data ($\mathcal{T}(n_{s}\in[16,128],n_{a}=5)$), including Gated Delta Net (GDN), Gated Slot Attention (GSA), Mamba2, and RWKV-7. Error decreases polynomially with context length up to 20k tokens, yielding an almost straight line when plotted against the logarithm of context length. Consequently, AnyMDP serves as a long-context benchmark for procedural memory, whereas prior benchmarks such as Needle-in-the-Haystack primarily probe episodic memory.
D.2 OmniRL achieves automatic trade-off between exploration and exploitation
@@
Previous studies have noted that in-context reinforcement learning (ICRL) can automatically balance exploration and exploitation. This phenomenon has been theoretically linked to posterior sampling. In Figure˜15, we illustrate the entropy of the decision-making process as a function of steps within the context. When compared to Figure˜5, we observe that the decrease in loss ($\mathcal{L}_{t}$) is primarily driven by the reduction in the entropy of the policy. Specifically, the agent initially assigns equal probabilities to all actions, reflecting an exploratory phase. As more contextual information accumulates, the agent gradually converges its choices, thereby transitioning towards exploitation. This empirical finding suggests that imitating an optimal policy (oracle) is sufficient to achieve an automatic balance between exploration and exploitation.
D.3 Additional Evaluation on Gymnasium
Figure˜16 and Figure˜17 demonstrate OmniRL’s online-RL, offline-RL, and imitation learning capabilities toward diverse unseen tasks.
@@ D.3 Additional Evaluation on Gymnasium
D.4 Memory states in ICRL implicitly encode the task structure
ICRL with linear attention captures all the information required to solve the environment in its memories ($\phi_{t}$). We perform a comprehensive t-SNE analysis to examine how these memories transform across different tasks during Online-RL evaluation. As shown in Figure˜18, the clustering patterns confirm the distinct task distributions of Gym, Darkroom, and AnyMDP. Notably, Darkroom and Gym clusters are predominantly located in the top-left region, while AnyMDP occupies a broader spatial area, reflecting its greater diversity. This spatial differentiation emphasizes AnyMDP’s unique characteristics and highlights OmniRL’s strong generalization ability across diverse tasks.
D.5 Comparison with Pre-trained LLMs
@@
You are playing the Frozen Lake game. The environment is a 4x4 grid where you need to maximize the success rate by reaching the goal (+1) without falling into holes (-1). You can move in four directions: left, down, right, and up (represented as 0, 1, 2, 3 respectively). You will receive the current state and need to provide the optimal action based on your learning. When asked for the optimal action, your response must be an integer ranging from 0 to 3, and no other context is permitted. There are two kind of request types: 1.integer: the integer is the current state, and you need to provide the optimal action. 2.list: The list contains one or more tuples, where each tuple contains the last state, action taken, reward received, and next state. To save time, you don’t need to respond when receiving a list. You will play the game multiple times. A game ends when the reward is -1 or 1, try to get a higher success rate. Note: I am asking you to play this game, not to find a coding solution or method. You will be provided with a conversation history. The latest prompt is the current state, and others are the list of sequential environment feedback history in tuple type. Each tuple contains four values, the first one is state, the second one is action, the third one is reward and the fourth one is next state. Your response must be an integer from 0 to 3 during the entire chat. If you find the last state is equal to the next state, your policy in the last state can’t be this action. If you find the reward in the tuple is -1, your policy in the last state can’t be this action. You need to get to the goal as soon as possible.
[…]
There is a game with the following basic description and rules: Frozen Lake involves crossing a frozen lake from the start to the goal without falling into any holes by walking over the frozen lake. The player may not always move in the intended direction due to the slippery nature of the frozen lake. The game starts with the player at location [0,0] of the frozen lake grid world, with the goal located at the far extent of the world, for example, [3,3] for the 4x4 environment. Holes in the ice are distributed in set locations when using a pre-determined map or in random locations when a random map is generated. The player makes moves until they reach the goal or fall into a hole. The lake is slippery, so the player may move perpendicular to the intended direction sometimes. If the intended direction is to the left, the actual move may be to the left, up, or down, with the corresponding probability distribution: P(move left) = 1/3, P(move up) = 1/3, P(move down) = 1/3. If the intended direction is to the right, the actual move may be to the right, up, or down, with the corresponding probability distribution: P(move right) = 1/3, P(move up) = 1/3, P(move down) = 1/3. If the intended direction is up, the actual move may be up, left, or right, with the corresponding probability distribution: P(move up) = 1/3, P(move left) = 1/3, P(move right) = 1/3. If the intended direction is down, the actual move may be down, left, or right, with the corresponding probability distribution: P(move down) = 1/3, P(move left) = 1/3, P(move right) = 1/3. You are given a 4x4 map where: S represents the start. F represents the frozen surface that can be walked on. H represents a hole; falling into it will return the player to the start. G represents the goal. The map is as follows: The first row from left to right is "SFFF". The second row from left to right is "FHFH". The third row from left to right is "FFFH". The fourth row from left to right is "HFFG". Please determine the optimal policy that maximizes the success rate of safely reaching the goal from the start. The optimal policy is the intended direction at each map location, where actions 0, 1, 2, and 3 represent moving left, down, right, and up, respectively. Note: You are not required to write code to solve this problem; instead, directly provide the optimal policy.
[…]
@@
In this particular environment, you should focus more on strategies to avoid encountering pitfalls rather than concentrating on reaching the goal. A reevaluation of the decision-making process in this context may yield more effective outcomes.
[…]
@@
The optimal policy for each state (row, column) is as follows, where each cell lists the intended direction (action) that maximizes the probability of reaching the goal: - **(0,0)**: Down (1) - **(0,1)**: Down (1) - **(0,2)**: Down (1) - **(0,3)**: Up (3) - **(1,0)**: Down (1) - **(1,2)**: Down (1) - **(2,0)**: Right (2) - **(2,1)**: Down (1) - **(2,2)**: Down (1) - **(3,1)**: Right (2) - **(3,2)**: Right (2) **Explanation**: - **Top row (0,0 to 0,3)**: Prioritize moving down to avoid the risky rightward path toward holes. - **Middle rows (1,0 and 1,2)**: Move down to progress toward safer lower rows. - **Lower rows (2,0 to 3,2)**: Navigate rightward toward the goal, avoiding holes. For example, (3,1) and (3,2) move right to reach the goal directly. - **(0,3)**: Move up (intended) to bias movement leftward (safer) and avoid falling into the hole at (1,3).
[…]
The optimal policy, revised to prioritize avoiding holes, is as follows: - **(0,0)**: Down (1) *Avoids (0,1)’s risks by favoring transitions to (1,0) or (0,1) without direct hole exposure.* - **(0,1)**: Right (2) *Moves toward (0,2) or (1,1) (hole). While risky, all actions here have 1/3 hole risk; right progresses toward safer (0,2).* - **(0,2)**: Right (2) *Directs to (0,3) instead of down to (1,2), which is high-risk.* - **(0,3)**: Up (3) *Biases movement left (to (0,2)) and avoids the hole at (1,3).* - **(1,0)**: Left (0) *Avoids 1/3 chance of moving right into (1,1) (hole). Transitions to (0,0) or (2,0) safely.* - **(1,2)**: Right (2) *Despite 1/3 hole risk, it allows progress to (2,2) or (0,2). Left (action 0) has equal risk but less progress.* - **(2,0)**: Up (3) *Avoids 1/3 chance of moving down into (3,0) (hole). Routes to (1,0) or (2,1).* - **(2,1)**: Down (1) *Safely moves to (3,1) (frozen), progressing toward the goal.* - **(2,2)**: Down (1) *Reaches (3,2), a safe step toward the goal.* - **(3,1)**: Right (2) *Directly progresses to (3,2), then (3,3) (goal).* - **(3,2)**: Right (2) *Final step to the goal (3,3).* ### Key Adjustments from Original Policy: 1. **(1,0)**: Changed to **left** (0) to avoid 1/3 chance of falling into (1,1). 2. **(1,2)**: Changed to **right** (2) instead of down, reducing hole risk from 2/3 to 1/3. 3. **(2,0)**: Changed to **up** (3) to avoid 1/3 chance of entering (3,0). 4. **(0,1)**: **Right** (2) minimizes repeated exposure to (1,1) compared to down. ### Path Safety Focus: - Avoids actions with >1/3 hole probability (e.g., down from (1,2)). - Prioritizes paths with lower immediate risk, even if longer (e.g., routing through (2,0) → (2,1) → (3,1) → goal).
[…]
@@
We also investigate whether a well-pretrained LLM can naturally solve the decision-making tasks investigated in this paper. To circumvent the lack of common sense in AnyMDP tasks, we primarily conducted tests in the FrozenLake task with DeepSeek-R1[77] in two modes:
@@
    Similar to the evaluation of standard ICRL, we do not provide the agent with the map. Instead, we report only the state ID and reward of the agent. The initial prompts used to initiate the evaluation are shown in Figure˜19.
@@
    We initially provide the global map to the DeepSeekR1DeepSeek-R1 and then commence the interaction. In this mode, the LLM can leverage the global map to make decisions. The prompts are shown in Figure˜20 and Figure˜21
As shown in Table˜8 and Figure˜24 (results excecuted by following the responses of DeepSeek-R1 (version 2025/03) in Figure˜22 and Figure˜23), LLM agents are only able to solve the FrozenLake (non-slippery) environment when provided with a global map. Without access to a global map, we conducted extensive interactions between LLM agents and the environment, running up to $500$ episodes ($100,000$ steps). Despite these efforts, the agents failed to solve even the non-slippery variant of the task, achieving scores that were comparable to those of a random policy.